# Lightweight Prompt Biasing for Contextualized End-to-End ASR Systems

### v1

@@ III-A Multi-task Learning for Contextual Biasing
TABLE I: Examples from models trained With/Without task tokens. The biasing list includes coffee, milk, chocolate.
System Task Tokens Output
Reference N/A So I’ve added four toffee almond milk hot cocoas what else?
Recognition Without coffee coffee milk hot cocoa no no no no no
Recognition With So I’ve added 4 toffee almond milk hot cocos what else?
Table I presents outputs from models trained with and without task tokens. The absence of task tokens leads the model to overfit to the prompt, resulting in repeated or irrelevant words (e.g., coffee) and reduced transcription accuracy. In contrast, incorporating task tokens enables the model to correctly identify and utilize relevant entities from the prompt while preserving accurate recognition of the remaining audio content. This example highlights that merely supplying prompt information is insufficient; without a multi-task framework and explicit task tokens, the model is prone to hallucinations and transcription errors. The proposed multi-task approach with task tokens allows the model to effectively differentiate between biasing and non-biasing scenarios, leveraging contextual information only when appropriate. This design enhances transcription accuracy and facilitates the integration of contextual biasing without introducing architectural complexity.
@@
1:  $hi$: index of the <hit> token
2:  $H\leftarrow\text{Encoder}(\text{audio})$Compute encoder embedding}
@@ III-B Entity Filtering for Biasing
Previous studies on CLAS have shown that attention-based neural biasing methods encounter difficulties when processing large biasing lists [13]. Our own experiments, as reflected in rows C1–C3 of Table II, corroborate this finding: large biasing lists introduce substantial noise, which diminishes the model’s ability to accurately focus on relevant entities. Our analysis reveals a clear correlation between the precision of the biasing list and the effectiveness of contextual biasing—the more targeted and relevant the biasing list, the greater the improvement in biasing performance. This observation motivates the introduction of an entity filtering mechanism to systematically eliminate irrelevant candidates from large biasing lists.
@@
As depicted in Figure 2, we adopt a progressive entity filtering strategy during decoding, utilizing the multi-task model to identify and exclude irrelevant entities from the biasing list. This process substantially reduces the size of the candidate list; for example, initial lists containing up to 2,000 entities can be efficiently narrowed to 10–20 highly relevant candidates. By eliminating distractors, the filtering mechanism enhances the model’s ability to focus on pertinent entities, thereby improving recognition accuracy. The detailed steps of this entity filtering approach are provided in Algorithm 1. Although the method requires evaluating each entity in the biasing list, it is highly amenable to parallelization and can be executed with minimal latency on modern hardware accelerators such as GPUs.
TABLE II: Performance on in-house domain dataset with various biasing lists.
@@ IV-A Training Details
We fine-tune our model from a pre-trained Transformer backbone, eliminating the need for training from scratch. The vocabulary is augmented with special tokens (<sop>, <hit>, and <miss>) to support the proposed multi-task learning framework. The training dataset comprises both biasing and non-biasing samples, with biasing samples constituting 65% of the total. For each sample, entities in the biasing list are randomly selected from either the reference transcription or a pool of negative phrases drawn from diverse text corpora, with each entity limited to a maximum of five words. Task tokens are assigned based on the presence of these entities in the audio. To manage memory consumption, each training sample contains fewer than 20 entities in the biasing list. The model is trained on approximately 5,400 hours of anonymized paired audio-text data using the AdamW optimizer [34] and a linear decay learning rate schedule, with a peak learning rate of $2.24\times 10^{-4}$ and a brief warmup phase.
@@ IV-B Evaluation Settings
We conducted a comprehensive evaluation of the proposed Prompt Biasing approach using an in-house domain-specific dataset covering 9 popular domains and comprising approximately 570,000 words. Each utterance was annotated by human experts to identify entities of interest. To measure recognition performance on these entities, we utilized the Entity Word Error Rate (EWER), which specifically quantifies errors within labeled entities.
To systematically investigate the effect of biasing list size, we constructed three types of biasing lists for each utterance:
@@ IV-C1 Contextual Biasing Performance
Rows C1–C3 in Table II present the performance of the Prompt Biasing model evaluated with three different types of biasing lists on our in-house domain dataset. First of all, when an ideal ground truth list (i.e. Exact list) is provided, the Prompt Biasing model achieves an impressive EWER of 1.8%, significantly outperforming its baseline model (i.e. A1 in Table II) by 76.8% relatively. This demonstrates the effectiveness of our approach in leveraging contextual information to enhance the entity recognition accuracy. In addition to the strong EWER performance, the Prompt Biasing model achieves a competitive WER of 4.60%, slightly outperforming the baseline. This indicated that the model effectively leverages contextual information to improve entity recognition while maintaining overall transcription quality. In addition to evaluating the Prompt Biasing model with the ideal Exact biasing list, we also assess its performance using two more realistic, noisy biasing lists: Small and Large. These lists better reflect practical scenarios where the biasing list may not be entirely accurate or precise. The Small list contains approximately 50 entities, while the Large list includes around 1800 entities. As shown in Table II (rows C2 and C3), the Prompt Biasing model achieves strong EWERs of 4.16% and 5.61% with the Small and Large lists, respectively, along with competitive WERs of 4.54% and 4.80%. These results demonstrate that the model maintains effective contextual biasing and robust general transcription accuracy even when provided with imperfect biasing lists.
@@ IV-C2 Impact of Entity Filtering
Given the observed benefits of providing a precise biasing list, we further evaluate the impact of our entity filtering strategy detailed in Algorithm 1. As shown in rows D1 and D2 of Table II, applying entity filtering leads to a notable improvement in EWER for both the Small and Large noisy biasing lists, compared to the standard prompt biasing model (rows C2 and C3). Specifically, EWER decreases from 4.16% to 4.04% for the Small list and from 5.61% to 4.95% for the Large list. These results demonstrate that entity filtering effectively enhances biasing performance, particularly when handling large and noisy biasing lists. Additionally, general transcription performance, as measured by WER, is also slightly improved—from 4.54% to 4.46% for the Small list and from 4.80% to 4.71% for the Large list—indicating that the entity filtering strategy not only improves entity recognition but also helps maintain overall transcription quality.
@@ IV-C3 Comparison to Shallow Fusion
Shallow fusion is a widely adopted technique for enhancing contextual biasing performance in ASR systems. As shown in Table II, we compare our proposed Prompt Biasing model with entity filtering (rows D1 and D2) against the baseline model with shallow fusion (rows B1 and B2) across different biasing list sizes. Our approach consistently outperforms shallow fusion, achieving a relative reduction in EWER of 30.7% for the Small biasing list (D1 vs. B1) and 18.0% for the Large biasing list (D2 vs. B2). These results highlight the effectiveness of Prompt Biasing with entity filtering in leveraging contextual information to improve entity recognition accuracy. Additionally, our method maintains competitive overall transcription quality, with WERs of 4.46% and 4.71% for the Small and Large lists, respectively, compared to 4.67% and 4.74% for shallow fusion. This demonstrates that our approach not only enhances entity recognition but also preserves general ASR performance, even with large and noisy biasing lists.
@@ IV-C4 Robustness Against Noise
Table III presents the performance of the models on a 7.6M-word in-house dataset under both standard (no biasing list) and noisy biasing list conditions. The standard setting corresponds to conventional ASR without contextual information. Compared to the baseline, the Prompt Biasing model exhibits only a marginal increase in WER of 0.04% in the absence of contextual information, demonstrating minimal impact on general ASR performance. Furthermore, when evaluated with a noisy biasing list, the Prompt Biasing model shows a negligible WER degradation of 0.06%, indicating strong robustness to irrelevant or inaccurate contextual information.
TABLE III: Performance on in-house dataset for robustness benchmark.

### v2

@@ III-A Multi-task Learning for Contextual Biasing
TABLE I: WERs of models trained With/Without task tokens on a food ordering dataset with/without biasing list.
Model trained
with Task Tokens
w/o biasing list w/ biasing list
No 10.11 19.21
Yes 9.70 9.46
TABLE II: Examples from models trained With/Without task tokens. The biasing list includes coffee, milk, chocolate.
System Result
Reference
So I’ve added four toffee almond milk
hot cocoas what else?
Model Without
Task Tokens
coffee coffee milk
hot cocoa no no no no no
Model With
Task Tokens
So I’ve added 4 toffee almond milk
hot cocos what else?
Table I summarizes initial results from an early investigation using a food ordering dataset, comparing models trained with and without the inclusion of specialized task tokens. The models were assessed under two settings: with a biasing list and without one (i.e., standard ASR). The findings reveal that models trained without task tokens experience a notable increase in WER when a biasing list is used, indicating an inability to effectively utilize contextual information from the biasing list. Conversely, models trained with task tokens achieve a marked reduction in WER under both conditions, demonstrating that these tokens are essential for leveraging contextual cues and focusing on relevant entities from the biasing list. In Table II, we present representative results obtained from the two models discussed above. The absence of task tokens leads the model to overfit to the prompt, resulting in repeated or irrelevant words (e.g., coffee, no ) and reduced transcription accuracy. In contrast, incorporating task tokens enables the model to correctly identify and utilize relevant entities from the prompt (e.g., milk) while preserving accurate recognition of the remaining audio content. This example highlights that merely supplying prompt information is insufficient; without a multi-task framework and explicit task tokens, the model is prone to hallucinations and transcription errors. The proposed multi-task approach with task tokens allows the model to effectively differentiate between biasing and non-biasing scenarios, leveraging contextual information only when appropriate. This design enhances transcription accuracy and facilitates the integration of contextual biasing without introducing architectural complexity.
@@
1:hihi: index of the <hit> token
[…]
@@ III-B Entity Filtering for Biasing
Previous studies on CLAS have shown that attention-based neural biasing methods encounter difficulties when processing large biasing lists [13]. Our own experiments, as reflected in rows C1–C3 of Table III, corroborate this finding: large biasing lists introduce substantial noise, which diminishes the model’s ability to accurately focus on relevant entities. Our analysis reveals a clear correlation between the precision of the biasing list and the effectiveness of contextual biasing—the more targeted and relevant the biasing list, the greater the improvement in biasing performance. This observation motivates the introduction of an entity filtering mechanism to systematically eliminate irrelevant candidates from large biasing lists.
@@
As depicted in Figure 2, we adopt a progressive entity filtering strategy during decoding, utilizing the multi-task model to identify and exclude irrelevant entities from the biasing list. This process substantially reduces the size of the candidate list; for example, initial lists containing up to 2,000 entities can be efficiently narrowed to 10–20 highly relevant candidates. By eliminating distractors, the filtering mechanism enhances the model’s ability to focus on pertinent entities, thereby improving recognition accuracy. The detailed steps of this entity filtering approach are provided in Algorithm 1. Although the method necessitates evaluating each entity in the biasing list, the additional computational overhead remains minimal. This efficiency is achieved by utilizing a shared encoder, eliminating the need for repeated computations. Furthermore, only a single forward pass through the relatively small decoder (approximately 100 million parameters) is required for each biasing entity, which can be efficiently batch-processed on modern hardware accelerators with minimal impact on latency.
TABLE III: Performance on in-house domain dataset with various biasing lists.
@@ IV-A Training Details
We fine-tune our model from a pre-trained Transformer backbone, eliminating the need for training from scratch. The vocabulary is augmented with special tokens (<sop>, <hit>, and <miss>) to support the proposed multi-task learning framework. The training dataset comprises both biasing and non-biasing samples, with biasing samples constituting 65% of the total. For each sample, entities in the biasing list are randomly selected from either the reference transcription or a pool of negative phrases drawn from diverse text corpora, with each entity limited to a maximum of five words. Task tokens are assigned based on the presence of these entities in the audio. To manage memory consumption, each training sample contains fewer than 20 entities in the biasing list. The model was trained on roughly 5,400 hours of anonymized paired audio-text data sourced from various domains, such as voice assistants, conversational speech, and dictation tasks, ensuring a diverse dataset that reflects real-world usage. Model parameters are optimized using the AdamW optimizer [34] with a linear decay learning rate scheduler that features a peak learning rate of $2.24\times 10^{-4}$ and includes a short warmup period.
TABLE IV: Number of unique entities in each domain dataset
Domain Banking Healthcare Enhanced Medicine
#Entity 973 1625 3143
Domain Gaming Insurance Law Enforcement
#Entity 1948 1296 1148
Domain Basic Medicine Vehicles Science Technology
#Entity 2633 1910 1584
@@ IV-B Evaluation Settings
We conducted a comprehensive evaluation of the proposed Prompt Biasing approach using an in-house domain-specific dataset covering 9 popular domains and comprising approximately 570,000 words. Each utterance was annotated by human experts to identify entities of interest. Table IV presents the number of unique entities identified in each domain-specific datasets. To measure recognition performance on these entities, we utilized the Entity Word Error Rate (EWER), which specifically quantifies errors within labeled entities. To systematically investigate the effect of biasing list size, we constructed three types of biasing lists for each utterance:
@@ IV-C1 Contextual Biasing Performance
Rows C1–C3 in Table III present the performance of the Prompt Biasing model evaluated with three different types of biasing lists on our in-house domain dataset. First of all, when an ideal ground truth list (i.e. Exact list) is provided, the Prompt Biasing model achieves an impressive EWER of 1.8%, significantly outperforming its baseline model (i.e. A1 in Table III) by 76.8% relatively. This demonstrates the effectiveness of our approach in leveraging contextual information to enhance the entity recognition accuracy. In addition to the strong EWER performance, the Prompt Biasing model achieves a competitive WER of 4.60%, slightly outperforming the baseline. This indicated that the model effectively leverages contextual information to improve entity recognition while maintaining overall transcription quality. In addition to evaluating the Prompt Biasing model with the ideal Exact biasing list, we also assess its performance using two more realistic, noisy biasing lists: Small and Large. These lists better reflect practical scenarios where the biasing list may not be entirely accurate or precise. The Small list contains approximately 50 entities, while the Large list includes around 1800 entities. As shown in Table III (rows C2 and C3), the Prompt Biasing model achieves strong EWERs of 4.16% and 5.61% with the Small and Large lists, respectively, along with competitive WERs of 4.54% and 4.80%. These results demonstrate that the model maintains effective contextual biasing and robust general transcription accuracy even when provided with imperfect biasing lists.
@@ IV-C2 Impact of Entity Filtering
Given the observed benefits of providing a precise biasing list, we further evaluate the impact of our entity filtering strategy detailed in Algorithm 1. As shown in rows D1 and D2 of Table III, applying entity filtering leads to a notable improvement in EWER for both the Small and Large noisy biasing lists, compared to the standard prompt biasing model (rows C2 and C3). Specifically, EWER decreases from 4.16% to 4.04% for the Small list and from 5.61% to 4.95% for the Large list. These results demonstrate that entity filtering effectively enhances biasing performance, particularly when handling large and noisy biasing lists. Additionally, general transcription performance, as measured by WER, is also slightly improved—from 4.54% to 4.46% for the Small list and from 4.80% to 4.71% for the Large list—indicating that the entity filtering strategy not only improves entity recognition but also helps maintain overall transcription quality.
@@ IV-C3 Comparison to Shallow Fusion
Shallow fusion is a widely adopted technique for enhancing contextual biasing performance in ASR systems. As shown in Table III, we compare our proposed Prompt Biasing model with entity filtering (rows D1 and D2) against the baseline model with shallow fusion (rows B1 and B2) across different biasing list sizes. Our approach consistently outperforms shallow fusion, achieving a relative reduction in EWER of 30.7% for the Small biasing list (D1 vs. B1) and 18.0% for the Large biasing list (D2 vs. B2). These results highlight the effectiveness of Prompt Biasing with entity filtering in leveraging contextual information to improve entity recognition accuracy. Additionally, our method maintains competitive overall transcription quality, with WERs of 4.46% and 4.71% for the Small and Large lists, respectively, compared to 4.67% and 4.74% for shallow fusion. This demonstrates that our approach not only enhances entity recognition but also preserves general ASR performance, even with large and noisy biasing lists.
@@ IV-C4 Robustness Against Noise
Table V presents the performance of the models on a 7.6M-word in-house dataset under both standard (no biasing list) and noisy biasing list conditions. The standard setting corresponds to conventional ASR without contextual information. Compared to the baseline, the Prompt Biasing model exhibits only a marginal increase in WER of 0.04% in the absence of contextual information, demonstrating minimal impact on general ASR performance. Furthermore, when evaluated with a noisy biasing list, the Prompt Biasing model shows a negligible WER degradation of 0.06%, indicating strong robustness to irrelevant or inaccurate contextual information.
TABLE V: Performance on in-house dataset for robustness benchmark.